OpenStudy (sarahc):

Find the derivative of f(x) = 6 divided by x at x = -2. 6 3 divided by 3 negative 3 divided by 2 -6

OpenStudy (anonymous):

f(x) = 6x^(-1) f'(x) = -6/x^(2) f'(2) = ?

OpenStudy (anonymous):

f'(-2)=?*

OpenStudy (solomonzelman):

10 minutes since, and no response. What is the problem, sarahc? If you don't understand ask, don't just sit there :)

OpenStudy (sarahc):

C?

OpenStudy (sarahc):

@ineptAtMath @SolomonZelman

OpenStudy (solomonzelman):

Yes